The points z1,z2,z3,z4 are veritces of parallelogram only if?

Grade:9

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

if they represent vertices of a parallelogram then
midpoint of diogonal coincides.
\\=>{(z_1+z_3) \over 2}={(z_2+z_4) \over 2} \\=>z_1+z_3=z_2+z_4

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Ptol
5 years ago
Z1 +z3 =z2 +z4 is the correct answer  of this question.I think u must have got it ,hope it helps